OpenStudy (anonymous):

y=-2x+490 Describe how you would graph this line using the slope-intercept method. Be sure to write using complete sentences.

OpenStudy (anonymous):

@uri

OpenStudy (uri):

@dan815 he's best in these so i won't take his place.

OpenStudy (anonymous):

@dan815 helpppppp

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

First off you need to find the x and y coordinates

OpenStudy (anonymous):

oooh

OpenStudy (dan815):

okay so, first when x=0 what is y

OpenStudy (dan815):

y=-2x+490 when x=0 the equation becomes y=-2*(0) + 490 =490 so we have this point (0,490) graph this first

OpenStudy (dan815):